OpenStudy (anonymous):

Please HELP!!! fan and medal!! How do i brush my teeth? a. with a toothbrush b. eat ice-cream c. pull the teeth out d. brick

OpenStudy (anonymous):

a. with a toothbrush

OpenStudy (anonymous):

A

OpenStudy (anonymous):

i think D? can someone work it out for me

OpenStudy (invinceble):

suck a d!ck

OpenStudy (antonio_xx2):

lol...A

OpenStudy (invinceble):

common sence ,stupid a

OpenStudy (anonymous):

sorry that was not a answer @invinceble maybe you need to get glasses

OpenStudy (anonymous):

is English not your native language?

OpenStudy (anonymous):

no this is MATH!

OpenStudy (antonio_xx2):

this is math:/

OpenStudy (anonymous):

get the tooth brush and go in and out lol i know you guys are thinking wrong

OpenStudy (anonymous):

See look il show you the equation to work it out

OpenStudy (bibby):

hopefully on a live cam site

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

first 2\[\le \subseteq2341\int\limits_{\sqrt[^{\tan^{-1} }]{?}}^{?}\]

OpenStudy (anonymous):

hmmmm first you need to know the trajectory of the angle of the tooth brush and the speed your brushing.

OpenStudy (anonymous):

and you get brick please help

OpenStudy (anonymous):

Teeth+tooth brush + angle = Tree i am very confused

OpenStudy (anonymous):

maybe you need to Graph it?

OpenStudy (anonymous):

if you clean your teeth with a brick you will eventually break your teeth or scrape it flat. ice cream is not considered as cleaning because your trying to clean the sugar, food, and etc off your teeth. Well if you pull your teeth out you can get serious infection like gingivitis. If you use a tooth brush you can clean it obviously better than all of the other choises.

OpenStudy (anonymous):

Thank you fried rice You have showed me that the Correct answer is icecream! Fan and medal for you!

OpenStudy (anonymous):

teeth + rectum + evangelism + ego = tree

OpenStudy (antonio_xx2):

:/

OpenStudy (anonymous):

omg everyone is dumb its ice cream... vanilla works the best but chocolate works too